OpenStudy (anonymous):

Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can do this 10 times. Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Phillip can receive.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

is the total number of points on a project is \(100.\) ?

OpenStudy (anonymous):

Im assuming so and I have to find the maximum he can make if he change things

OpenStudy (anonymous):

this is worded bad

OpenStudy (jdoe0001):

hmm thought it was 100 myself as well though it's really unclear it's 100 the only clear part is that whatever amount was previous making changes gives him the next term, which is 2/10 of the total and that he can do that 10 times only may not amount to 100 though

OpenStudy (jdoe0001):

hint: \(\begin{array}{cccccclllllll} term&&a_1&a_2&a_3&... \\\hline\\ value&&75&75+\frac{2}{10}\ of \ {\color{purple}{ 75}}& \left( 75+\frac{2}{10}\textit{of 75}\right)+\frac{2}{10}\ of {\color{purple}{ \left( 75+\frac{2}{10}\textit{of 75}\right) }}&... \end{array}\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

okay. it says create the formula so...

OpenStudy (jdoe0001):

actually "the missing points back" are not 25... so .... hmm

OpenStudy (jdoe0001):

are not 75 rather

OpenStudy (jdoe0001):

so if we assume is 100 he's got 75 already the "missing points" are 25 so each change makes up for 2/10 of that

OpenStudy (anonymous):

okay. do i use this Sn =(1-r^n)/1-r formula?

OpenStudy (mathmath333):

the formula is correct, but the values we are assuming should be correct.

OpenStudy (anonymous):

im confused now. Like where are we

OpenStudy (mathmath333):

for instance the value this is giving is in million http://www.wolframalpha.com/input/?i=75*%5Cfrac%7B1-%2825*2%2F10%29%5E%2810%29%29%7D%7B1-25*2%2F10%7D%3D

OpenStudy (jdoe0001):

hmmm I read it as \(\large \begin{array}{lllllll} \begin{array}{llll} term\\changes \end{array}&value \\\hline\\ a_1&75\\ a_2&75+\frac{2}{10}\ of \ {\color{purple}{ 25}}\\ a_3&75+\frac{4}{10}\ of \ {\color{purple}{ 25}}\\ a_4&75+\frac{6}{10}\ of \ {\color{purple}{ 25}}\\ a_5&75+\frac{8}{10}\ of \ {\color{purple}{ 25}}\\ a_6&75+\frac{10}{10}\ of \ {\color{purple}{ 25}}\\ a_7&75+\frac{11}{10}\ of \ {\color{purple}{ 25}}\\ ...&... \end{array}\)

OpenStudy (mathmath333):

but thats not geometric series right

OpenStudy (anonymous):

thats not

OpenStudy (anonymous):

ive been trying to do this for an hour

OpenStudy (mathmath333):

the question says to create the formula and it doesn't specifically asking the value

OpenStudy (anonymous):

but i don't know what to write

OpenStudy (mathmath333):

may be just write this ,not sure though. \(\large \color{black}{\begin{align} \sum_{i=1}^{10}\text{nth term}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

idk thank for trying